At 425 K, this reaction has a Kc value of 0.0116.
2X (g)+2Y (g) **arrows left and right** 2Z (g)
calculate Kp at 425 K.
Kp=?

Prist CalcuatorPerlodc Table uion 4 of 16 At 425 K, this reaction has a Ke value of 0.0116 Calouate K at 425 K about us careers pertners prkacy pa
Answer #1

use the following expression: Kp = Kc(RT)dn

dn is the difference of the coefficients of the products and reactants, which in this case:
dn = (2) - (2+2) = -2

Using R = 0.0821 L atm/mol K and T = 425 K let's calculate Kp:

Kp = 0.0116 / (0.0821*425)2
Kp = 9.53x10-6

Hope this helps